# Weighted blow-up desingularization of plane curves

### v1

@@ Abstract.
Given a singular hypersurface in a regular 2-dimensional scheme essentially of finite type over a field, we construct an embedded resolution of singularities by tame Artin stacks. This applies in particular to plane curves. The key tool is Hironaka’s characteristic polyhedron from which we determine a suitable center for a stack-theoretic weighted blow-up. In particular, we show that the order strictly decreases at every point lying above the center. In contrast to earlier work by the authors, we do not have to restrict to perfect base fields and we deduce an inductive argument, despite the fact that higher dimensional tangent spaces arise, by taking torus actions and equivariant centers into account. Hence, we work with weighted blow-ups of tame Artin stacks and we do not have to pass to multi-weighted blow-ups of Deligne–Mumford stacks.
@@ 1. Introduction
This paper is a continuation of [AQS], where we revisited the resolution of singularities of plane curves using stack-theoretic blow-ups. In that paper we considered a singular curve $C\subset S$ embedded in a smooth surface over a field $K$ which we assumed perfect. We showed in [AQS, Theorem 1.1] that there is a natural weighted blow-up $S^{\prime}\to S$, where the proper transform $C^{\prime}\subset S^{\prime}$ of $C$ has lower order. The surface $S^{\prime}$ is, however, in general an Artin stack of the form $[\mathcal{W}^{\prime}/\mathbb{G}_{m}]$, where the action has possibly inseparable stabilizers, and $\mathcal{W}^{\prime}$ is a 3-dimensional scheme. In order to be able to use induction, in [AQS, Theorem 1.2] we replaced $S^{\prime}$ by another stack-theoretic blowup $S^{\prime\prime}\to S$, which is a Deligne–Mumford stack, allowing us to repeat the process.
In this paper we take a more direct approach, using the weighted blowup $S^{\prime}=[\mathcal{W}^{\prime}/\mathbb{G}_{m}]$ itself. In order to be able to repeat blowing up $S^{\prime}$, we expand the generality where we allow higher dimensional schemes, but require them to possess a torus actions. The precise setup is as follows:
@@
    Let $\bar{J}=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$ be the associated reduced center. Let $B\to\mathcal{W}$ be the associated degeneration to the normal cone of $\bar{J}$, with a natural open subscheme $\mathcal{W}^{\prime}:=B_{+}\subset B$ admitting a $T^{\prime}=T\times\mathbb{G}_{m}$-action lifting the $T$-action on $\mathcal{W}$. Then the action of $T^{\prime}$ on $\mathcal{W}^{\prime}$ has finite stabilizers, and the morphism $[\mathcal{W}^{\prime}/T^{\prime}]\to[\mathcal{W}/T]$ is the weighted blow-up along $\bar{J}$, which is a proper and birational morphism of tame Artin stacks.
@@
    Let $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}$ be the proper transform of $\mathcal{Z}$. Then the order of $\mathcal{Z}^{\prime}$ at every point lying over $\mathcal{Q}$ is strictly smaller than $a_{1}$. In particular, the invariant defined in Part (1) of the center associated to $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}$ at every point of $\mathcal{W}^{\prime}$ lying over $\mathcal{Q}$ is strictly smaller than $(a_{1},a_{2})$.
Iterating $\mathcal{W}^{\prime}\to\mathcal{W}$, one obtains in finitely many steps $\mathcal{W}^{(m)}\to\mathcal{W}^{(m-1)}$, with the reduction $\mathcal{Z}^{(m)}_{\operatorname{red}}\subset\mathcal{W}^{(m)}$ nonsingular. Taking quotients $[\mathcal{Z}^{(m)}/T^{(m)}]_{\operatorname{red}}\subset[\mathcal{W}^{(m)}/T^{(m)}]$ we obtain an embedded resolution of singularities of $[\mathcal{Z}/T]\subset[\mathcal{W}/T]$ by tame Artin stacks.
@@
An application of Bergh’s destackification [B] provides a scheme-theoretic embedded resolution of $\mathcal{Z}\subset\mathcal{W}$ and of $[\mathcal{Z}/T]\subset[\mathcal{W}/T]$, see [AQS, Remark 5.6].
Rather than proving the result directly on higher dimensional schemes, we pass to the localization at the generic point of $\mathcal{Q}$ to reduce to the case of a singular hypersurface $C\subset S$ on a regular 2-dimensional scheme. This approach requires the following upgrade of [AQS, Theorem 1.1] to the case where $S$ is essentially of finite type over a field:
@@
    The order of the proper transform of $C$ at every point of the blow-up lying over $q$ is strictly smaller than $a_{1}$.
@@ 2. A brief introduction to Hironaka’s characteristic polyhedron
On the other hand, if $K$ is imperfect and if $[K^{2}(\lambda):K^{2}]=2$, i.e., $\lambda$ is not a square, then $v=2$ is not solvable.
@@ 3. Invariant, parameters, and center
Let us address the situation if (2.1) does not hold. Then $\operatorname{in}_{\mathfrak{m}}(f):=f\mod\mathfrak{m}^{\nu+1}\in\operatorname{gr}_{\mathfrak{m}}(\mathcal{O})$ cannot be expressed as polynomial in a single variable. For readers familiar with Hironaka’s notion of the directrix (for example, see [CJS, Definitions 2.8 and 2.26]), the latter means that the dimension of the directrix is zero. Due to a result of Hironaka [H2, Theorem 2] which was generalized by Mizutani [M] (see also [CJS, Theorem 3.14]) blowing up the closed point $q$ will lead in this situation to a strict decrease of the order at points lying above $q$.
@@
    For resolution of $1$-dimensional schemes the local perspective via valuations is enough. In general, for instance in Theorem 1.1, one has to describe centers globally, as for example in [ATW, Section 2.2] via valuative $\mathbb{Q}$-ideals.
@@
    $x_{2}=y_{2}$ : By construction, we may vary $x_{2}$ for the given $x_{1}$ as long as we have $\mathfrak{m}=(x_{1},\widetilde{x_{2}})$ for the different choice. The analogous statement is true for $y_{2}$. Since $\mathbb{P}(\mathfrak{m}/\mathfrak{m}^{2})$ has at least 3 points, we may assume without loss of generality that $x_{2}=y_{2}$. Further, this provides $v_{J_{y}}(x_{2})=v_{J_{x}}(x_{2})=\frac{1}{a_{2}}$.
@@ 4. Weighted blow-ups and proper transform
Let $B=\operatorname{Spec}_{S}(\tilde{R})$ be the spectrum relative to the scheme $S$ of the quasi-coherent sheaf of $\mathcal{O}_{S}$-algebras $\tilde{R}$. Besides from its structure morphism $B\to S$, $B$ is also equipped with a morphism $\varphi\colon B\to\mathbb{A}_{k}^{1}=\operatorname{Spec}k[s]$. Using the latter $B$ becomes the degeneration of $S$ to the weighted normal cone of $\bar{J}$. Using Equation (4.1), we have:
@@
Let $B=\operatorname{Spec}_{S}(\tilde{R})$ be as in Construction 4.2. The weighted blow-up of $S$ along $\bar{J}$ is defined as the Artin stack
@@
Consider an expansion of $f$ as in Equation (3.2), $f=f_{J,1}(x_{1},x_{2})+\widetilde{h}(x_{1},x_{2})$. Recall that $v_{\bar{J}}(f)=v_{\bar{J}}(f_{J,1})=\ell<v_{\bar{J}}(\widetilde{h})$ where $\ell\in\mathbb{Z}_{+}$ is the integer introduced in Definition 4.1. Using the presentation of $\tilde{R}$ given in Equation (4.1), the pullback of $f$ to $B$ is
@@
The argument in the Section 5 below guarantees that its order along the exceptional locus $s=0$ is $<2p$. This follows from the fact that the term $y^{\prime 2p}+x^{\prime m}y^{\prime p}z^{m+p}+x^{\prime 2m}z^{p+1}$ is not of the form $g^{2p}$ and that $x^{\prime}$ or $y^{\prime}$ is invertible (and also remember that $z$ is invertible).
@@ 5. Proofs
The order drops: Proof of Theorem 1.3(3).
Since the order is functorial for smooth morphisms, showing that the order drops on $B_{+}$ implies that it drops on the quotient ${\operatorname{Bl}}_{\bar{J}}(S)=[B_{+}/\mathbb{G}_{m}]$. Note that $\nu_{\mathfrak{m}}(f)\leq\nu^{\log}_{\mathfrak{m}}(f)$, where the logarithmic order $\nu^{\log}_{\mathfrak{m}}(f)$ is defined with respect to the exceptional divisor $\{s=0\}$: it is the order where $s\neq 0$ and otherwise it is the order of the restriction to $\{s=0\}$, namely the order of $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})$.
When $s\neq 0$ we simply are computing the order of $f$ away from $q=\{x_{1}=x_{2}=0\}$. We note that near $q$ this order is $<\nu=a_{1}$ since $f$ is not a pure $\nu$-th power.
When $s=0$ we are computing the order of $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})$ away from $\{x_{1}^{\prime}=x_{2}^{\prime}=0\}$. But once again, $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})=x_{1}^{\prime\nu}+\cdots$ is not a pure $\nu$-th power, see Corollary 3.6. Thus its order of vanishing at any point on this locus is $<\nu=a_{1}$, as needed. ∎
@@
    Firstly, letting $k^{\operatorname{sep}}$ denote the separable closure of $k$, we claim $U$ is stable under $T(k^{\operatorname{sep}})$. Indeed, $J$ remains the unique center of maximal invariant $(a_{1},a_{2})$ that is admissible for $C\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})\subset S\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})$ at $q$. On the other hand, for every $t\in T(k^{\operatorname{sep}})$, $t\cdot J$ is also a center on $S\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})$ that satisfies the same properties, so by the uniqueness of $J$, $t\cdot J=J$. Thus, $J=t\cdot J$ extends to $t\cdot U\subset\mathcal{Q}$, so $t\cdot U\subset U$.
@@
    Secondly, we claim that in fact $U$ is stable under $T$. To do so, we instead show that $F=\mathcal{Q}\smallsetminus U$ is stable under $T$, or equivalently, that the preimage of $F$ under the $T$-action $\alpha\colon T\times\mathcal{Q}\to\mathcal{Q}$ contains $T\times F$. By (i), $\alpha^{-1}(F)\supset\{t\}\times F$ for every $t\in T(k^{\operatorname{sep}})$. Since $T$ is integral and geometrically reduced, the points in $T(k^{\operatorname{sep}})$ are dense in $T$. Thus, after taking closures, we deduce $\alpha^{-1}(F)\supset T\times F$ as desired.
@@
We now prove the remaining parts of Theorem 1.1. Following the notation introduced in Section 4, consider the weighted blow-up of $[B_{+}/\mathbb{G}_{m}]\to\mathcal{W}$ along $\bar{J}=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$. Since $B_{+}\subset B=\operatorname{Spec}(\tilde{R})$ with $\mathcal{O}_{\mathcal{W}}\subset\tilde{R}\subset\mathcal{O}_{\mathcal{W}}[t,t^{-1}]$, the co-action $\mathcal{O}_{\mathcal{W}}\to\mathcal{O}_{\mathcal{W}}\otimes\mathcal{O}_{T}$ naturally lifts to a co-action $\tilde{R}\to\tilde{R}\otimes\mathcal{O}_{T}$ by sending $t\mapsto t\otimes 1$. In this way, the $T$-action on $\mathcal{W}$ lifts to an action of $T^{\prime}=T\times\mathbb{G}_{m}$ on $B_{+}$, inducing a morphism $[B_{+}/T^{\prime}]\to[\mathcal{W}/T]$ fitting in a cartesian diagram:
${{[B_{+}/\mathbb{G}_{m}]}}$${{[B_{+}/T^{\prime}]}}$${\mathcal{W}}$${{[\mathcal{W}/T]}}$
Since $\mathcal{W}\to[\mathcal{W}/T]$ is a smooth cover, and $[B_{+}/\mathbb{G}_{m}]\to\mathcal{W}$ is proper and birational by Lemma 4.4, $[B_{+}/T^{\prime}]\to[\mathcal{W}/T]$ is also proper and birational. In order to show $[B_{+}/T^{\prime}]$ has finite stabilizers, it suffices to show its relative inertia over $[\mathcal{W}/T]$ is finite, since by hypothesis $[\mathcal{W}/T]$ has finite stabilizers. Indeed, that is the case, as the relative inertia of $[B_{+}/T^{\prime}]$ over $[\mathcal{W}/T]$ is finite after base change to the smooth cover $\mathcal{W}\to[\mathcal{W}/T]$.
Finally, by part (3) of Theorem 1.3, the order of $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}:=B_{+}$ at every point lying over $q$ is strictly smaller than $a_{1}$. Since every point lying over $\mathcal{Q}$ is in the same $T^{\prime}$-orbit as points lying over $q$, we deduce the same conclusion for points in $\mathcal{Z}^{\prime}$ lying over $\mathcal{Q}$. ∎

### v2

# Torus actions, weighted blow-ups, and desingularization of plane curves

@@ Abstract.
Given a singular hypersurface in a regular 2-dimensional scheme essentially of finite type over a field, we construct an embedded resolution of singularities by weighted blow-ups. This differs from our earlier work which required multi-weighted blow-ups. We deduce an inductive argument, despite the fact that higher dimensional tangent spaces arise, by taking torus actions and equivariant centers into account. In addition, we do not have to restrict to perfect base fields.
@@ 1. Introduction
This paper is a continuation of [AQS], where we revisited the resolution of singularities of plane curves using stack-theoretic blow-ups. The two papers broadly share the same motivation, which we now recall.
We are concerned with resolution of singularities in positive characteristics. This is in general a challenging topic, however, as in [AQS], we focus on the case of plane curves, known since the early days of the subject, beginning with Newton.
A new method of resolution in characteristic 0, using stack-theoretic weighted blow-ups, was introduced in [ATW, Mc], and one must wonder if this method has implications in positive characteristics, how it relates to existing methods, and what possible hurdles it encounters. Specifically, colleagues have asked us to elaborate the first case, of plane curves.
In the earlier paper [AQS] we thus considered a singular curve $C\subset S$ embedded in a smooth surface over a field $K$ which we assumed perfect. We showed in [AQS, Theorem 1.1] that there is a natural weighted blow-up $S^{\prime}\to S$, where the proper transform $C^{\prime}\subset S^{\prime}$ of $C$ has lower order. This stood in contrast with earlier methods, as order reduction traditionally necessitates multiple steps. [^1] The exception is the method of Puiseux exponents of unibranch complex curve singularities. It is worth mentioning that the local transformations there agree with those determined by the weighted blow-up method, see the first sections of [Cu, K]. The surface $S^{\prime}$ is, however, in general an Artin stack of the form $\mathcal{W}^{\prime}/\mathbb{G}_{m}]$, where the action has possibly inseparable stabilizers, and $\mathcal{W}^{\prime}$ is a 3-dimensional scheme. In order to be able to use induction, in [AQS, Theorem 1.2] we replaced $S^{\prime}$ by another stack-theoretic blowup $S^{\prime\prime}\to S$, which is a Deligne–Mumford stack, where the order is still reduced after blowing up, allowing us to repeat the process.
This left us with the question whether we can complete a treatment of weighted resolution of plane curves directly, using the weighted blowup $S^{\prime}=[\mathcal{W}^{\prime}/\mathbb{G}_{m}]$ itself. This is what we achieve in this paper. This has the benefit that the logarithmic order of the singularity drops at each step, which implies that the order drops as well.
In order to be able to repeat blowing up $S^{\prime}$, we expand the generality where we allow higher dimensional schemes, but require them to possess torus actions. In contrast with [AQS], our methods require us to leave the perfect field case. The precise setup is as follows:
@@
    Let $\bar{J}=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$ be the associated reduced center. Let $B\to\mathcal{W}$ be the associated degeneration to the normal cone of $\bar{J}$, with a natural open subscheme $\mathcal{W}^{\prime}:=B_{+}\subset B$ admitting a $T^{\prime}=T\times\mathbb{G}_{m}$-action lifting the $T$-action on $\mathcal{W}$. Then the action of $T^{\prime}$ on $\mathcal{W}^{\prime}$ has finite stabilizers, and the morphism $\mathcal{W}^{\prime}/T^{\prime}]\to[\mathcal{W}/T]$ is the weighted blow-up along $\bar{J}$, which is a proper and birational morphism of tame Artin stacks.
@@
    Let $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}$ be the proper transform of $\mathcal{Z}$. Then the logarithmic order of $\mathcal{Z}^{\prime}$ with respect to the exceptional divisor at every point lying over $\mathcal{Q}$ is strictly smaller than $a_{1}$. In particular, the invariant defined in Part (1) of the center associated to $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}$ at every point of $\mathcal{W}^{\prime}$ lying over $\mathcal{Q}$ is strictly smaller than $(a_{1},a_{2})$.
Iterating $\mathcal{W}^{\prime}\to\mathcal{W}$, one obtains in finitely many steps $\mathcal{W}^{(m)}\to\mathcal{W}^{(m-1)}$, with the reduction $\mathcal{Z}^{(m)}_{\operatorname{red}}\subset\mathcal{W}^{(m)}$ nonsingular. Taking quotients $\mathcal{Z}^{(m)}/T^{(m)}]_{\operatorname{red}}\subset[\mathcal{W}^{(m)}/T^{(m)}]$ we obtain an embedded resolution of singularities of $\mathcal{Z}/T]\subset[\mathcal{W}/T]$ by tame Artin stacks.
@@
An application of Bergh’s destackification [B] provides a scheme-theoretic embedded resolution of $\mathcal{Z}\subset\mathcal{W}$ and of $\mathcal{Z}/T]\subset[\mathcal{W}/T]$, see [AQS, Remark 5.6].
Rather than proving the result directly on higher dimensional schemes, we pass to the localization at the generic point of $\mathcal{Q}$ to reduce to the case of a singular hypersurface $C\subset S$ on a regular 2-dimensional scheme. This approach requires the following upgrade of [AQS, Theorem 1.1] to the case where $S$ is essentially of finite type over an arbitrary field:
@@
    The logarithmic order of the proper transform of $C$ with respect to the exceptional divisor at every point of the blow-up lying over $q$ is strictly smaller than $a_{1}$.
@@ 2. A brief introduction to Hironaka’s characteristic polyhedron
On the other hand, if $K$ is imperfect and if $K^{2}(\lambda):K^{2}]=2$, i.e., $\lambda$ is not a square, then $v=2$ is not solvable.
@@ 3. Invariant, parameters, and center
Let us address the situation if (2.1) does not hold. Then $\operatorname{in}_{\mathfrak{m}}(f):=f\mod\mathfrak{m}^{\nu+1}\in\operatorname{gr}_{\mathfrak{m}}(\mathcal{O})$ cannot be expressed as polynomial in a single variable. For readers familiar with Hironaka’s notion of the directrix (for example, see [CJS, Definitions 2.8 and 2.26]), the latter means that the dimension of the directrix is zero. Due to a result of Hironaka [H2, Theorem 2] which was generalized by Mizutani [Mi] (see also [CJS, Theorem 3.14]) blowing up the closed point $q$ will lead in this situation to a strict decrease of the order at points lying above $q$.
@@
    For resolution of 11-dimensional schemes the local perspective via valuations is enough. In general, for instance in Theorem 1.1, one has to describe centers globally, as for example in [ATW, Section 2.2] via valuative $\mathbb{Q}$-ideals.
@@
[…]
@@ 4. Weighted blow-ups and proper transform
[…]
@@
[…]
@@
[…]
@@
The argument in the Section 5 below guarantees that its order along the exceptional locus $s=0$ is <2p<2p. This follows from the fact that the term $y^{\prime 2p}+x^{\prime m}y^{\prime p}z^{m+p}+x^{\prime 2m}z^{p+1}$ is not of the form $g^{2p}$ and that $x^{\prime}$ or $y^{\prime}$ is invertible (and also remember that $z$ is invertible).
@@ 5. Proofs
The logarithmic order drops: Proof of Theorem 1.3(3).
Since the logarithmic order with respect to the exceptional divisor $\{s=0\}$ is functorial for smooth morphisms, showing that the logarithmic order drops on $B_{+}$ implies that it drops on the quotient ${\operatorname{Bl}}_{\bar{J}}(S)=[B_{+}/\mathbb{G}_{m}]$. Note that $\nu_{\mathfrak{m}}(f)\leq\nu^{\log}_{\mathfrak{m}}(f)$, where the logarithmic order $\nu^{\log}_{\mathfrak{m}}(f)$ is defined: it is the order where $s\neq 0$ and otherwise it is the order of the restriction to $\{s=0\}$, namely the order of $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})$.
When $s\neq 0$ we simply are computing the order of $f$ away from $q=\{x_{1}=x_{2}=0\}$. We note that near $q$ this order is $\nu=a_{1}$ since $f$ is not a pure $\nu$-th power.
When $s=0$ we are computing the order of $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})$ away from $\{x_{1}^{\prime}=x_{2}^{\prime}=0\}$. But once again, $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})=x_{1}^{\prime\nu}+\cdots$ is not a pure $\nu$-th power, see Corollary 3.6. Thus its order of vanishing at any point on this locus is $\nu=a_{1}$, as needed. ∎
@@
    Letting $k^{\operatorname{sep}}$ denote the separable closure of $k$, we claim $U$ is stable under $T(k^{\operatorname{sep}})$. Indeed, $J$ remains the unique center of maximal invariant $(a_{1},a_{2})$ that is admissible for $C\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})\subset S\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})$ at $q$. On the other hand, for every $t\in T(k^{\operatorname{sep}})$, $t\cdot J$ is also a center on $S\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})$ that satisfies the same properties, so by the uniqueness of $J$, $t\cdot J=J$. Thus, $J=t\cdot J$ extends to $t\cdot U\subset\mathcal{Q}$, so $t\cdot U\subset U$.
@@
    We claim that in fact $U$ is stable under $T$. To do so, we instead show that $F=\mathcal{Q}\smallsetminus U$ is stable under $T$, or equivalently, that the preimage of $F$ under the $T$-action $\alpha\colon T\times\mathcal{Q}\to\mathcal{Q}$ contains $T\times F$. By (i), $\alpha^{-1}(F)\supset\{t\}\times F$ for every $t\in T(k^{\operatorname{sep}})$. Since $T$ is integral and geometrically reduced, the points in $T(k^{\operatorname{sep}})$ are dense in $T$. Thus, after taking closures, we deduce $\alpha^{-1}(F)\supset T\times F$ as desired.
@@
[…]
$B_{+}/\mathbb{G}_{m}]$$B_{+}/T^{\prime}]$$\mathcal{W}$$\mathcal{W}/T]$}}
Since $\mathcal{W}\to[\mathcal{W}/T]$ is a smooth cover, and $B_{+}/\mathbb{G}_{m}]\to\mathcal{W}$ is proper and birational by Lemma 4.4, $B_{+}/T^{\prime}]\to[\mathcal{W}/T]$ is also proper and birational. In order to show $B_{+}/T^{\prime}]$ has finite stabilizers, it suffices to show its relative inertia over $\mathcal{W}/T]$ is finite, since by hypothesis $\mathcal{W}/T]$ has finite stabilizers. Indeed, that is the case, as the relative inertia of $B_{+}/T^{\prime}]$ over $\mathcal{W}/T]$ is finite after base change to the smooth cover $\mathcal{W}\to[\mathcal{W}/T]$.
Finally, by Theorem 1.3(3), the logarithmic order of $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}:=B_{+}$ at every point lying over $q$ is strictly smaller than $a_{1}$. Since every point lying over $\mathcal{Q}$ is in the same $T^{\prime}$-orbit as points lying over $q$, we deduce the same conclusion for points in $\mathcal{Z}^{\prime}$ lying over $\mathcal{Q}$. ∎